OpenStudy (anonymous):

What is the slope of a line that passes through the point (-5,3) and is parallel to a line that passes through (2,13) and (-4,-11)

OpenStudy (anonymous):

x+5/-6=y-3/-24 -4x-20+3=y y=-4x-17 the slope is -4